INTERNATIONAL CONFERENCE ON MATHEMATICAL ANALYSIS AND ITS APPLICATIONS, 2000 (ICMAA2000)

January 17 - 21, 2000 National Sun Yat-sen University, Kaohsiung, Taiwan, R.O.C.

Second Announcement June 22, 1999

The aim of ICMAA2000 is to bring together mathematicians working in Abstract and Applied Analysis to enhance the interaction among areas of research. Currently, we have arranged:

\underline{Keynote speakers}: Louis de Branges (Purdue), Fon-Che Liu (Academia Sinica), Gilles Pisier (Texas A&M and Paris VI), Ralph Rockafellar (Washington) and Mau-Hsiang Shih (Chung Yuan Christian U.).

\underline{Invited speakers}: Y. Abramovich (IUPUI), F. Altomare (U. Bari), T. Ando (Hokusei Gakuen University), W. Arendt (U. Ulm), G. Barsegian (Nat’l Acad. Sci. Armenia), L. G. Brown (Purdue), P. G. Casazza (Missouri-Columbia), J. Chabrowski (Queensland), K. J. Chen (Tsing Hua U.), S. S. Cheng (Tsing Hua U.), M. D. Choi (Toronto), C. H. Chu (London), C. P. Chu (Soochow U.), C. C. Cowen (Purdue), B. D. Craven (Melbourne), G. Ding (Nankai), Y. F. Fang (Cheng Kung U.), F. Giannessi (U. Pisa), Y. He (Chinese Academy Sci.), F. Hiai (Tohoku), T. Hida (Meijo), M. J. Huang (Tsing Hua U.), K. M. Hui (Academia Sinica), S. Igari (Tohoku), C. R. Johnson (William and Mary), N. J. Kalton (Missouri-Columbia), W. A. Kirk (Iowa), S. Koshi (Hokkaido Inst. Tech.), C. C. Kuo (Fu Jen Catholic U.), I. Laine (U. Joensuu), A. T. Lau (Alberta) , D. Leung (Singapore), C. K. Li (William and Mary), Borluh Lin (Iowa), P. K. Lin (Memphis), S. Lu (Beijing Normal U.), S. L. Lu (Tung-Hau U.), M. Mimura (Hiroshima U.), K. Mizukami (Hiroshima U.), F. Moricz (Szeged), R. Nagel (U. Tuebingen), D. Nualart (U. Barcelona), W. Oettli (Mannheim), I. V. Ostrovskii (Turkey U.), S. Park (Seoul), G. Pedersen (Copenhagen) , B. Russo (UC Irvine), Z. Shi (Harbin), A. H. Siddiqi (KFUTM, Saudi Arabia), A. P. Sigh (U. Jammu), L. Streit (U. Bielefeld), W. Takahashi (Tokyo Inst. of Tech.), B. S. Tam (Tamkang U.), K. Tanaka (Niigata Inst. Tech.), T. Tanaka (Hirosaki U.), S. L. Troyanski (Sofia), S. Y. Tzeng (Changhau U. Ed.), G. X. Z. Yuan (Queensland), S. Yotsutani (Ryukoku), Zhongrui Shi (Harbin), A. S. Ustunel (ENST, France), P. Q. Vu (Ohio), H. C. Wang (Tsing Hua U.), J. N. Wang (Cheng Kung U.), T. Wang (Harbin), Y. Wang (Chinese Academy Sci.), C. Wu (Harbin), S. Y. Wu (Cheng Kung U.).

\underline{Keynote/Invited Speakers to be confirmed}: P. Cohen (Stanford), M. Yor (Pairs VI)

\underline{Mathematicians who may give a talk or participate in ICMAA2000}: R. Ahmad (India) Y. Aihara (Japan), H. Akca (Saudi Arabia), Y. Alber (Israel), I. Andrianov (Ukraine), R. Aron (USA), R. Aulaskari (Finland), I. Beg (Kuwait), W. Bergweiler (Germany), S. L. Blyumin (Russia), J. A. van Casteren (Belgium), S. Chen (China), Q. Cheng (Australia), D. Dinh (Vietnam), B. Efim (Russia), D. Fan (USA), J. Gao (USA), C. G. Hu (China), P. C. Hu (China), X. H. Hua (China) Z. Huang (China), N. Igor (Russia), N. Ishimura (Japan), P. Jain (India), K. Jarosz (USA), S. B. Joshi (India), A. Karlovich (Ukraine), M. Kisaka (Japan), H. Kriete (Germany), N. Komuro (Japan), S. Leonid (Ukraine), B. S. Lee (Korea), L. Li (Australia), P. Li (China), F. Lust-Piquard (France), A. McIntosh (Australia), L. Maligranda (Sweden), V. M. Manuilov (Russia), E. Merzbach (Israel), T. W. Ng (England), T. Nishishiraho (Japan), Ho Ou (USA), J. Y. Park (Korea), J. Puerto (Spain), B. Ramazan (Romania), B. Randrianantoanina (USA), N. Randrianantoanina (USA), Sadallah (Algeria), N. Sakakibara (Japan), S. Saitoh (Japan), N. Sergei (Russia), M. Shapiro (USA), D. Shoikhet (Israel), A. G. Siskakis (Greece), T. Y. Tam (USA), K. Taira (Japan), T. Takenaka (Japan), M. Taniguchi (Japan), T. H. Thao (Vietnam), K. Tohge (Japan), T. Toshiharu (Japan), L. M. Tovar (Mexico), V. K. Tuan (Kuwait), K. Victor (Israel), B. Walther (Sweden), C. T. Wu (Germany), J. Wu (China), S. J. Wu (China), H. Wulan (China), Y. Z. Yin (China), C. J. Zhang (China), G. Y. Zhang (China), K. Zhang (Australia), Z. Ziegler (Israel).

Anybody interested in the conference is warmly invited to attend and to give a talk. Please contact any one of the following organizing committee members for further information:

Banach Spaces of Analytic Functions Mark C. Ho (hom@math.nsysu.edu.tw)
Banach Space Theory  Borluh Lin (bllin@pop.math.uiowa.edu)
Convex Analysis  Jen-Chih Yao (yaojc@math.nsysu.edu.tw)
Fourier Analysis  Chang-Pao Chen (cpchen@math.nthu.edu.tw)
KKM and Fixed Point Theory  Lai-Jiu Lin (Maljlin@math.ncue.edu.tw)
Matrix Analysis  Mau-Hsiang Shih (mhshih@math.cycu.edu.tw)
Nonlinear Analysis  Hwai-Chiuan Wang (hwang@math.nthu.edu.tw)
Nonlinear PDEs  Jong-Shenq Guo (jsguo@math.ntnu.edu.tw)
Numerical Ranges  Pei-Yuan Wu (pywu@cc.nctu.edu.tw)
Operator Algebras  Ngai-Ching Wong (wong@math.nsysu.edu.tw)
Operator Semigroups  Sen-Yen Shaw (shaw@math.ncu.edu.tw)

and Evolution Equations

Optimization Theory  Hang-Chin Lai (hclai@csa500.isu.edu.tw)
Stochastic Analysis  Yuh-Jia Lee (yjlee@mail.ncku.edu.tw)
Value distribution theory  Chung-Chun Yang (mayang@uxmail.ust.hk)

and complex dynamics

Proceedings: The Proceedings of ICMAA2000 will be published as a special issue of the Taiwanese Journal of Mathematics (TJM). All speakers in the conference are encouraged to make contribution to it. The submitted papers will go through the normal refereeing process. Only those meeting the standards of the TJM will be accepted. The submission deadline is May 31, 2000 and the special issue is scheduled to appear in March, 2001. In preparing the manuscript, please follow the TJM Guidelines for Authors, which together with other information of the TJM can be found in its web site at http://www.math.nctu.edu.tw/tjm/. The manuscripts are to be submitted to one of the following editors: Yuh-Jia Lee (National Cheng Kung University), Borluh Lin (The University of Iowa), Sen-Yen Shaw (National Central University), Mau-Hsiang Shih (Chung-Yuan University), and Pei Yuan Wu (National Chiao Tung University).

Social Events: There will be two sight-seeing tours. One of them (tentatively on Wednesday) will take a half day to visit the city Kaohsiung, which includes a boat trip to look over the harbour. Another one (tentatively on Friday) will be a one day trip to the Kenting National Park which is located at the southern-most part of Taiwan island and is full of charming scene. The cost for each of them is about US$25 including meals.

Dinners, Lunches and Snacks: The Conference will provide lunch boxes to participants. There will be conference dinners/banquets. Together with coffee, tea and refreshments provided in between the lectures, all of them are free of charge to participants.

Venue: The Conference will be held mainly at National Sun Yat-sen University, Kaohsiung. Parts of the program may be held at National Cheng Kung University, Tainan, and I-Shou University, Kaohsiung.

Accommodations: The following hotels have promised to give participants of ICMAA2000 special discounts. Prices quoted below are after discount.

• Kingship Hotel (98, Chi-Hsien 3rd Road; Tel: 886-7-5313131): Single NT$1,100, Double (two beds) NT$1,960, For Three NT$ 2,100 and For Four NT$2,520 (including breakfast and service charges).

• Chinatrust Hotel (43, Tajen Road; Tel: 886-7-5217111): Single NT$1,660 and Double NT$1,860 (including breakfast, free airport transportation, and service charges).

• Howard Plaza Hotel (331, Chi-Hsien 1st Road; Tel: 886-7-2362323): Single/Double NT$2,000 (including breakfast, free airport transportation, free transportation between the University and the Hotel, and service charges).

Registration Fee: The Organizing Committee of ICMAA2000 is seeking supports from the Mathematics Development and Promotion Center, National Science Council of the Republic of China, the Ministry of Education of the Republic of China, the Mathematics Society of the Republic of China, National Sun Yat-sen University, National Cheng Kung University, I-Shou University, and other sources. These funds, however, are limited. There will be a small registration fee of US$100 (subject to change), and no charge for students. Please pay the registration fee in cash in US dollars at the registration desk in the Conference. People from local area may expect to pay the fee at a reduced rate (NT$600, subject to change) due to certain special support program. We encourage all participants to ask for support from their home universities or other institutions.

Travel Information:

• If you travel by air, please make sure that \underline{your final destination is Hsiaokang International Airport at Kaohsiung}.

The airport code of Kaohsiung Airport shown in your tickets will be KHH. There is another (much bigger) Chiang Kai-shek International Airport located at Taoyuan, which is 5 hours highway driving from Kaohsiung. We advise you to approach directly to the service desk of the hotel of your reservation upon getting out from the custom of Kaohsiung Airport. Shuttle services are provided by some hotels. In case the hotel you are staying does not provide this service or you arrive off office hours, you can take a taxi to the hotel or the university. You are supposed to pay by meter plus a fixed airport taxi service charge (NT$50). The cost for a single trip from Kaohsiung Airport to a hotel in downtown or to the University is about NT$300 + 50. Tips is not necessary. However, you have to prepare some local currency and, if you are not a Chinese speaker, a piece of paper in Chinese instruction with English subtitles, like the ones we will provide with the next (third) announcement, to tell the driver where you would like to go. You can also take a bus to Kaohsiung Train Station at the bus stop located right at the exit of the domestic airlines building which is conjuncted to the international airlines building in 5 minute walking distance. The fare is NT$25 per one way trip. Then see next paragraph.

• If you travel by train, you can take a taxi from Kaohsiung Train Station to the University for about NT$120 per single trip. Or you can take the No. 1 bus there as an alternative. You should get off from the bus at the University Tunnel Entrance bus stop. The fare is NT$10 (or NT$12 for air-conditioned buses).

• If you travel by car, exit at the Chung Cheng Road ramp on Highway 1. Then turn right at the first traffic light and you will be on Chung Cheng Road. Follow Chung Cheng Road until the end for about 20 minutes. Soon after you pass the Love River Bridge, make a right turn at the loop. You will be on Dagung Road then. Just a few block, you will be on the joint of a ‘T’ shape bridge. Make a left turn there. Drive straight and make a right turn at Lin Hai 2nd Road. After four traffic lights you will arrive at the University Tunnel Entrance (for bicycles and walkers only). Make a left turn and follow the ‘S’ sharp river side Lienhai road. About 5 minutes you will see the Main Entrance of the National Sun Yat-sen University. Tell the guards there you are attending ICMAA2000 and then you will know how to drive to the venue.

• If you plan to arrive after Monday, please note that we might have a one-day program held at Cheng Kung University or I-Shou University in other cities. These update will be shown in the Third Announcement.

Visas and other information: Citizens of some countries may enter Taiwan without visas at the Chiang Kai-shek or Hsiaokang International Airports. These include citizens from Australia, Austria, Belgium, Canada, France, Germany, Japan, Luxembourg, New Zealand, Portugal, Spain, Sweden, the Netherlands, the United Kingdom, and the United States. Visa-free entry allows stays of up to 14 days starting from midnight on the day of arrival. Please call your travel agent or the Taiwan, ROC, embassies, consulates or designated representative offices in your countries to make sure. Phone numbers of some of them are: Austria (61-3) 96508611, Belgium (32-2) 511-0687, Brazil (55-61) 3491218, Canada (1-416) 369-9030, Denmark (45-31) 197-511, Finland (358-0) 680-1216, France (33-1) 443-98820, Germany (49-30) 861-2754, Great Britain (44-71) 396-9152, Greece (30-1) 687-6750, Hong Kong (852) 2525-8315, Hungary (361) 366-2884, Ireland (353-1) 678-5413, Israel (972-3) 544-0250, Italy (39-6) 8841362, Japan (81-92) 473-6655, Korea (82-2) 399-276, Kuwait (965) 533-9988, Liberia Republic (231) 224-044, Mexico (525) 5207851, Netherlands (31-70) 346-9438, New Zealand (64-9) 303-3903, Poland (48-2) 630-8438, Portugal (351) 386-0617, Russia (750-3) 9563786, Saudi Arabia (966-1) 188-1900, South Africa (27-11) 339-8147, Singapore (65) 278-6511, Spain (34-1) 571-4729, Sweden (46-8) 728-8513, Switzerland (41-31) 382-2927, Thailand (66-2) 251-9393, Turkey (90-312) 436-7255, United States (1-202) 895-1800, and Vietnam (848) 8299-343. One may find other useful information from the web page of the Tourism Bureau, ROC at http://www.tbroc.gov.tw. A brief history of Taiwan and travel information are available there, too. The temperature of Kaohsiung in January is about 15–20 degrees C. No heavy clothing is in need.

Currency Exchange: The current exchange rate is about US$1=NT$33 (New Taiwan Dollars). It is best to make exchange at Kaohsiung Airport right after you pass the custom. You can also make exchange at banks in downtown or at the hotels. Almost no service charge or surplus applies when you do it in a bank or at the airport. You can change it back to US dollars or other currency at the airport again before you leave the country. Credit cards are popular here.

Further information: Please contact

Ngai-Ching Wong, Department of Applied Mathematics, National Sun Yat-sen University, Kaohsiung 80424, Taiwan, R.O.C.

Yuh-Jia Lee, Department of Mathematics, National Cheng Kung University, Tainan 701, Taiwan, R.O.C.

Borluh Lin, Department of Mathematics, The University of Iowa, Iowa City, IA 52242, U.S.A.

E-Mail: icmaa2000@math.nsysu.edu.tw Fax: 886-7-5253809

www-site: http://www.math.nsysu.edu.tw/u/icmaa2000, which is mirrored at
         http://www.math.uiow.edu/icmaa2000

In case you are interested to give a talk in ICMAA2000, please send us (and also the corresponding organizing committee member) a title and a short abstract at the latest by September 30, 1999. We ask you to understand that the number of lectures in the parallel sessions is limited, so that we may not be able to accommodate every proposed lecture in these sessions.

Please expect for the Third Announcement by December 1999.

Please complete and return the form in next page (please do it again even if you have already done so before since additional lodging information is needed), to, preferably by e-mail: icmaa2000@math.nsysu.edu.tw, or by Fax: 886-7-5253809, or via regular mail to

Ngai-Ching Wong  Yuh-Jia Lee  Borluh Lin
Department of Applied Mathematics Department of Mathematics  Department of Mathematics
National Sun Yat-sen University  National Cheng Kung University The University of Iowa
Kaohsiung 80424  Tainan 701  Iowa City, IA 52242
Taiwan, R.O.C.  Taiwan, R.O.C.  U.S.A.

You can also fill in the form at our www home page at

http://www.math.nsysu.edu.tw/u/icmaa2000, http://www.math.uiow.edu/icmaa2000.

Please note that we have to receive the title and abstract of your talk (if you are giving one in ICMAA2000) before September 30, 1999, and the lodging information (if you would like to ask the Conference to make reservation for you) before November 30, 1999. We cannot guarantee successful arrangement for talks and lodging after these dates.

Registration and lodging request form for ICMAA2000

□ I intend to participate in the International Conference on Mathematical Analysis

and its Applications, 2000, held at Kaohsiung, Taiwan, ROC, for Jan. 17–21, 2000. □ I propose to give a talk.

□ Title and abstract are attached.

□ Title and abstract will be submitted no later than September 30, 1999.

Name: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Institution: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Address: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ... . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Electronic mail: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Phone: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Fax: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Day of Arrival: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Day of Departure: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Accommodations:

□ I would like to arrange the lodging myself.

□ I would like to stay in the following Hotel and ask the Conference to make reservation for nights from the night of to the night of for my party of persons. I fill 1, 2, and 3 in the blanks below to indicate my preference in order.

Kingship Hotel

□ Single Room NT$1,100 □ Double Room NT$1,960 □ Room for Three NT$2,100 □ Room for four NT$2,520 Chinatrust Hotel (Free transportation between airport and hotel) □ Single Room NT$1,660 □ Double Room NT$1,860 Howard Plaza Hotel (Free transportation between airport, hotel and venue) □ Single Room NT$2,000 □ Double Room NT$2,000

Registration Fees: □ US$100 for overseas guests (payable at the registration desk during the Conference) □ 0 for students/special guests □ NT$600 for local people (partially supported by sponsors) Remarks: